Check whether T(x1,x2,x3)=(x1+x2/x3,x3) define linear transformation from R3 to R2


1
Expert's answer
2021-06-14T17:57:34-0400

Let us check whether "T(x_1,x_2,x_3)=(x_1+\frac{x_2}{x_3},x_3)" define a linear transformation from "\R^3" to "\R^2."

Taking into account that

"T(x_1,x_2,x_3)+T(y_1,y_2,y_3)=(x_1+\frac{x_2}{x_3},x_3)+(y_1+\frac{y_2}{y_3},y_3)=(x_1+y_1+\frac{x_2}{x_3}+\frac{y_2}{y_3},x_3+y_3)"

and "T(x_1+y_1,x_2+y_2,x_3+y_3)=(x_1+y_1+\frac{x_2+y_2}{x_3+y_3},x_3+y_3)\ne T(x_1,x_2,x_3)+T(y_1,y_2,y_3),"

we conclude that "T(x_1,x_2,x_3)=(x_1+\frac{x_2}{x_3},x_3)" does not define a linear transformation from "\R^3" to "\R^2."